# Mastering Multi-Drone Volleyball through Hierarchical Co-Self-Play Reinforcement Learning

### v3

@@ C.1 High-Level Strategy
where $\mathrm{win\_or\_lose}_{j}=1$ if team $j$ wins, $-1$ if it loses, and 00 otherwise; $\mathrm{racket\_hit\_ball}_{j}=1$ if any racket of team $j$ contacts the ball at timestep $t$, and 00 otherwise.
@@ C.2 Training Result
The results of high‑level strategy training are shown in Fig. 10. In our experiments, we set the maximum population size to five, yielding five iterative training phases; each phase terminates when either a preset number of training steps is reached or a win‑rate threshold is met. The lower triangular portion of the matrix reports the win rate of each row policy against each column policy (i.e. the probability that the row policy prevails). Because the game is symmetric and zero‑sum, the upper‑triangular entries are complementary, such that each pair of win rates sums to one.
Figure 10: High-Level Strategy Training Result
@@ C.3 Hyperparameters
buffer length $64$ num minibatches $16$ PPO epochs $4$
@@
max population size $5$ win rate threshold $0.9$ max iteration steps $1.6384\times 10^{8}$
@@ D.1 Ablation Study on Low-Level Skill Refinement Schedule
Instead of jointly co-optimizing the high-level strategy with all low-level skills simultaneously, we co-optimize the high-level strategy with one low-level skill at a time (Algo. 1). We compare the two co-optimization methods by evaluating the win rates of the saved checkpoints against the high-level strategy and low-level skills obtained after Stage II, under identical settings. These settings include the reward structure, maximum number of iterations ($M=4000$), checkpoint saving interval ($N=1000$), and win rate threshold ($t=0.7$). Each iteration collects $32768$ sampled steps. All results are averaged over three random seeds. As shown in Fig. 11, jointly co-optimizing the high-level strategy with all low-level skills leads to near-zero win rates. In contrast, co-optimizing the high-level strategy with one low-level skill at a time can yield checkpoints that achieve win rates above 50%. Since MARL is inherently non-stationary, simultaneously updating all low-level skills destabilizes the learning process. In comparison, optimizing one low-level skill at a time reduces non-stationarity and enables more stable and effective learning.
@@
0:  Low-level skill set $\{\pi^{L}\}$, high-level policy $\pi^{H}$, number of max iterations $M$, checkpoint saving interval $N$, win rate threshold $t$
@@
4:     ${\pi^{L}}^{\prime}\leftarrow\pi^{L}$
5:     ${\pi^{H}}^{\prime}\leftarrow\pi^{H}$
6:     for $i=1$ to $M$ do
7:        Jointly optimize ${\pi^{L}}^{\prime}$ and ${\pi^{H}}^{\prime}$ through self-play against frozen $\{\pi^{L}\}$ and $\pi^{H}$
8:        if $i\bmod N=0$ then
9:           Save a checkpoint of ${\pi^{L}}^{\prime}$ and ${\pi^{H}}^{\prime}$
10:        end if
11:     end for
12:     Evaluate all saved checkpoints of ${\pi^{L}}^{\prime}$ and ${\pi^{H}}^{\prime}$ win rates against frozen $\{\pi^{L}\}$ and $\pi^{H}$
13:     Select the checkpoint of ${\pi^{L}}^{\prime}$ with the highest win rate, denoted as $\pi^{L}_{\text{best}}$
14:     if win rate of $\pi^{L}_{\text{best}}>t$ then
15:        Add $\pi^{L}_{\text{best}}$ to $\Pi_{\text{new}}$
16:     end if
17:  end for
18:  return  $\{\pi^{L}\}\cup\Pi_{\text{new}}$
Figure 11: Comparison of win rates under different co-optimization training schedules.
@@ D.2 Ablation Study on KL Divergence Penalty
As stated in Sec. 3.3, we incorporate a KL divergence penalty term into the low-level skill reward, in addition to the game-result-related reward shared with the high-level strategy. Here, we conduct a detailed ablation study on the KL divergence penalty to investigate its influence on training performance. As shown in Fig. 12, and following Algo. 1, we compare different settings of the KL penalty by evaluating the win rates of the saved checkpoints against the high-level strategy and low-level skills obtained after Stage II. We observe that a large KL penalty coefficient (e.g., $1\mathrm{e}{-2}$) causes the policy performance to collapse rapidly; conversely, without a KL penalty (coefficient = 0), the policy hovers only slightly above zero win rate. Only with a moderate coefficient (e.g., $1\mathrm{e}{-4}$) does the policy yield checkpoints achieving win rates above 50%. This is because, without a KL penalty, the reward signal for low-level strategies becomes sparse, significantly slowing down learning. On the other hand, an overly large KL penalty forces the policy to remain too close to the original one, which leads to accumulated error over time as the policy rigidly imitates previous behaviors without adaptation.
Figure 12: Comparison of win rates under different KL divergence coefficients.
@@ D.3 Ablation Study on High-Level Strategy Parameter Initialization
As the previous high-level strategy becomes invalid due to changes in the size of the low-level skill pool, it is necessary to retrain the high-level strategy to accommodate the augmented skill set. The initialization of this new high-level strategy is a critical factor influencing the overall performance in Stage III. We compare two different initialization methods for the high-level strategy. The first is a straightforward approach that trains the high-level strategy from scratch. In the second approach, we expand each actor head’s parameter vector to match the larger action space by copying over its original values. Any newly added dimensions are then initialized to the mean of that head’s original parameters. In both settings, we train the new high-level strategy against the frozen high-level strategy and low-level skills obtained after Stage II. As shown in Fig. 13, although training from scratch requires more time steps to converge, it ultimately achieves a higher win rate. In contrast, the parameter-averaging initialization method tends to converge to a suboptimal local solution.
Figure 13: Comparison of win rates under different high-level strategy parameter initializations.
@@ F.2 Nash-Averaging Evaluator
While head-to-head win rates fail to reflect absolute policy strength in non-transitive games, and Elo ratings [22] exhibit sensitivity to the composition of the policy population, Nash-averaging evaluator, introduced by Liu et al. [21], provides a robust alternative. This method selects a randomly sampled set of policies, then constructs a payoff matrix from pairwise match-ups, and finally computes a Nash equilibrium (NE). The resulting equilibrium weights yield a comparatively population-invariant measure of relative performance, mitigating biases induced by cyclic dominance. So we adopt this as an evaluation metric. The details are as follows.
@@
The meta policy is derived by computing the NE of the $10\times 10$ win rate matrix. Since the underlying game is zero-sum, win rates (probabilistic outcomes) must first be converted into zero-sum payoff values within the range $[-1,1]$, where one agent’s gain is the other’s loss. Specifically, the conversion follows:

### v4

@@ C.1 High-Level Strategy
where $\mathrm{win\_or\_lose}_{j}=1$ if team $j$ wins, 1-1 if it loses, and 0 otherwise; $\mathrm{racket\_hit\_ball}_{j}=1$ if any racket of team $j$ contacts the ball at timestep $t$, and 0 otherwise.
@@ C.2 Training Result
The results of high‑level strategy training are shown in Fig. 10. In our experiments, we set the maximum population size to five, yielding five iterative training phases; each phase terminates when either a preset number of training steps is reached or a win‑rate threshold is met. The lower triangular portion of the matrix reports the win rate of each row policy against each column policy (i.e., the probability that the row policy prevails). Because the game is symmetric and zero‑sum, the upper‑triangular entries are complementary, such that each pair of win rates sums to one.
Figure 10: Win-rate heatmap illustrating the evolution of high-level strategy training in Stage II.
@@ C.3 Hyperparameters
buffer length $64$ num minibatches $16$ PPO epochs 44
@@
max population size 55 win rate threshold $0.9$ max iteration steps $1.6384\times 10^{8}$
@@ D.1 Ablation Study on Low-Level Skill Refinement Schedule
Instead of jointly co-optimizing the high-level strategy with all low-level skills simultaneously, we co-optimize the high-level strategy with one low-level skill at a time (Algo. 1). We compare the two co-optimization methods by evaluating the win rates of the saved checkpoints against the high-level strategy and low-level skills obtained after Stage II, under identical settings. These settings include the reward structure, maximum number of iterations ($M=4000$), checkpoint saving interval ($N=1000$), and win rate threshold ($t=0.7$). Each iteration collects $32768$ sampled steps. All results are averaged over three random seeds. As shown in Fig. 11(a), jointly co-optimizing the high-level strategy with all low-level skills leads to near-zero win rates. In contrast, co-optimizing the high-level strategy with one low-level skill at a time can yield checkpoints that achieve win rates above 50%. Since MARL is inherently non-stationary, simultaneously updating all low-level skills destabilizes the learning process. In comparison, optimizing one low-level skill at a time reduces non-stationarity and enables more stable and effective learning.
@@
0: Low-level skill set $\{\pi^{L}\}$, high-level policy $\pi^{H}$, number of max iterations $M$, checkpoint saving interval NN, win rate threshold $\tau$
@@
4:  ${\pi_{k}^{L}}^{\prime}\leftarrow\pi_{k}^{L}$
5:  ${\pi_{k}^{H}}^{\prime}\leftarrow\pi^{H}$
6:  Form team $(\{\pi^{L}\}\setminus\{\pi_{k}^{L}\})\cup\{\pi^{L^{\prime}}_{k}\}$ with high-level $\pi^{H^{\prime}}_{k}$
7:  for $i=1$ to $M$ do
8:   Jointly optimize $\pi_{k}^{L}$$^{\prime}$ and $\pi_{k}^{H}$$^{\prime}$ through self-play against the frozen team $\{\pi^{L}\}$ and $\pi^{H}$
9:   if $i\bmod N=0$ then
10:    Save a checkpoint of $\pi_{k}^{L}$$^{\prime}$ and $\pi_{k}^{H}$$^{\prime}$
11:   end if
12:  end for
13:  For each $\pi^{L^{\prime}}_{k},\pi^{H^{\prime}}_{k}$ checkpoint, form team $(\{\pi^{L}\}\setminus\{\pi_{k}^{L}\})\cup\{\pi^{L^{\prime}}_{k}\}$ with high-level $\pi^{H^{\prime}}_{k}$ and evaluate the win rate against the frozen team $\{\pi^{L}\}$ and $\pi^{H}$.
14:  Select the checkpoint of $\pi_{k}^{L}$$^{\prime}$ with the highest win rate, denoted as $\pi^{L}_{k,\text{best}}$
15:  if win rate of $\pi^{L}_{k,\text{best}}>\tau$ then
16:   Append $\pi^{L}_{k,\text{best}}$ to $\Pi_{\text{new}}$
17:  end if
18:end for
19:return $\{\pi^{L}\}\cup\Pi_{\text{new}}$
(a) Comparison of win rates under different co-optimization training schedules.
(b) Comparison of win rates under different KL divergence coefficients.
(c) Comparison of win rates under different high-level strategy parameter initializations.
Figure 11: Ablation study results in Stage III.
@@ D.2 Ablation Study on KL Divergence Penalty
As stated in Sec. 3.3, we incorporate a KL divergence penalty term into the low-level skill reward, in addition to the game-result-related reward shared with the high-level strategy. Here, we conduct a detailed ablation study on the KL divergence penalty to investigate its influence on training performance. As shown in Fig. 11(b), and following Algo. 1, we compare different settings of the KL penalty by evaluating the win rates of the saved checkpoints against the high-level strategy and low-level skills obtained after Stage II. We observe that a large KL penalty coefficient (e.g., $1\mathrm{e}{-2}$) causes the policy performance to collapse rapidly; conversely, without a KL penalty (coefficient = 0), the policy hovers only slightly above zero win rate. Only with a moderate coefficient (e.g., $1\mathrm{e}{-4}$) does the policy yield checkpoints achieving win rates above 50%. This is because, without a KL penalty, the reward signal for low-level strategies becomes sparse, significantly slowing down learning. On the other hand, an overly large KL penalty forces the policy to remain too close to the original one, which leads to accumulated error over time as the policy rigidly imitates previous behaviors without adaptation.
@@ D.3 Ablation Study on High-Level Strategy Parameter Initialization
As the previous high-level strategy becomes invalid due to changes in the size of the low-level skill pool, it is necessary to retrain the high-level strategy to accommodate the augmented skill set. The initialization of this new high-level strategy is a critical factor influencing the overall performance in Stage III. We compare two different initialization methods for the high-level strategy. The first is a straightforward approach that trains the high-level strategy from scratch. In the second approach, we expand each actor head’s parameter vector to match the larger action space by copying over its original values. Any newly added dimensions are then initialized to the mean of that head’s original parameters. In both settings, we train the new high-level strategy with frozen augmented low-level skills against the frozen high-level strategy and low-level skills obtained after Stage II. As shown in Fig. 11(c), although training from scratch requires more time steps to converge, it ultimately achieves a higher win rate. In contrast, the parameter-averaging initialization method tends to converge to a suboptimal local solution.
@@ Appendix E Baselines
To ensure fairness, we provide all baseline agents with the full-state observations of opponent drones as those available to HCSP. Additionally, HCSP is trained independently without exposure to any of the baseline agents, guaranteeing an unbiased evaluation.
@@ F.2 Nash-Averaging Evaluator
While head-to-head win rates fail to reflect absolute policy strength in non-transitive games, and Elo ratings [22] exhibit sensitivity to the composition of the policy population, the Nash-averaging evaluator, introduced by Liu et al. [21], provides a robust alternative. This method selects a randomly sampled set of policies, then constructs a payoff matrix from pairwise match-ups, and finally computes a Nash equilibrium (NE). The resulting equilibrium weights yield a comparatively population-invariant measure of relative performance, mitigating biases induced by cyclic dominance. So we adopt this as an evaluation metric. The details are as follows.
@@
The meta policy is derived by computing the NE of the $10\times 10$ win rate matrix. Since the underlying game is zero-sum, win rates (probabilistic outcomes) must first be converted into zero-sum payoff values within the range [1,1][-1,1], where one agent’s gain is the other’s loss. Specifically, the conversion follows:
@@
Appendix G Preliminary Real-World Experiments
(a) Drone setup
(b) Serve
(c) Solo Bump
(d) Ball trajectory in Serve
(e) Height variation in Solo Bump
Figure 12: Real-world experiments.
Although the current method is not yet sufficient for full 3v3 zero-shot deployment, we incorporate key real-world factors in simulation. Our preliminary Sim2Real adaptations include:
  • We support two action space options. In addition to the 50Hz PRT, we also support 50Hz CTBR policy output followed by an 800Hz PID controller.
  • We add both domain randomization (initial states and restitution coefficient) and observation latency.
We further conduct preliminary real-world experiments to demonstrate the zero-shot Sim2Real potential of our approach on learned low-level skills, such as the Serve skill and a Solo Bump task, in which the drone is required to bump the ball repeatedly to showcase its agility and stability. As shown in Fig. 12(a), to stay within the quadrotor’s thrust margin, we mount a badminton racket (diameter 22cm, mass 60g) on top of the drone and use a small foam ball (diameter 4cm, mass 47g). Both the drone and the ball are tracked by a motion-capture system: the drone, modeled as a rigid body, receives position and orientation directly from mocap, while an Extended Kalman filter fuses these data with PX4 IMU readings to estimate velocity; the ball, treated as a point mass, also gets position from mocap, and a Kalman filter differentiates these measurements to recover velocity. The policies are deployed on the onboard Nvidia Orin processor. In zero-shot trials, the low-level skills such as Serve are executed successfully (Fig. 12(b) and 12(d)) and the Solo Bump policy achieves 29 consecutive bumps (Fig. 12(c) and 12(e)). These results provide promising evidence for the physical plausibility and transferability of our approach.